buss of mass 200kg travels at a uniform velocity 40m/s for distance of 2500 m before it come to rest. Calculate average deceleration.


1
Expert's answer
2021-03-16T11:35:38-0400

Deceleration can be determined from the provided information about braking distance and speed:


"a=\frac{v_f^2-v_i^2}{2d}=\frac{0^2-40^2}{2·2500}=0.32\text{ m/s}^2."